OpenStudy (18jonea):

What is the approximate molar mass of a molecular solute if 30 g of the solute in 100 g of water causes the solution to have a boiling point of 101°C? (Kb=0.512°C/m: Kf=1.86 °C/m; molar mass of water =18 g) 15 amu 300 amu 150 amu 30 amu

OpenStudy (18jonea):

@Cuanchi

OpenStudy (18jonea):

@aaronq

OpenStudy (18jonea):

@nincompoop

OpenStudy (aaronq):

Do you have any idea on how to use the information given to figure it out?

OpenStudy (18jonea):

not really

OpenStudy (aaronq):

You have to use the boiling point elevation equation, find the molality, then find the moles of solute, then finally (knowing it's mass) find the molar mass. Solve for m, molality: \(\sf \Delta T=m*K_b\) Solve for n, moles: \(\sf m=\dfrac{n}{Kg_{solvent}}\) Solve for molar mass: \(\sf n=\dfrac{mass}{Molar~mass}\)

OpenStudy (18jonea):

so how would i figure it out @aaronq

OpenStudy (aaronq):

i kinda already broke it down for you, solve for each one of the variables in the equations i wrote.

OpenStudy (18jonea):

15 @aaronq

OpenStudy (aaronq):

hm that's not correct, can you briefly post your calculations?

OpenStudy (18jonea):

then 150?

OpenStudy (18jonea):

@aaronq

OpenStudy (aaronq):

yup